Light of wavelength 2000A falls on an aluminium surface (work function of Al = 4.2ev ) to cause photoelectric emission. the stopping of potential in volt of the metal is??

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Aishwarya

According to Einstein's photoelectric equation

  hv = Φ0 + (KE)max = Φ0 + eVs, where Φis work function and Vs is stopping potential
  hv= hc/λ= 6.633 x 10-34 x 3 x 108 x/2000 x 10-10
     = 9.95 x 10-19 J = 9.95/1.6 eV= 6.22 eV
eVs= 6.22 - 4.2 = 2.02 eV
Vs= 2.02 volt

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 



Aiswarya Ram Gupta
13 years ago
thnx..